### v1

@@ 1 Introduction
There are various physical quantities used to measure the entanglement of quantum systems. Among them, entanglement entropy is a well-defined entanglement measure for pure states. However, it is not suitable as a measure for the entanglement of mixed states Horodecki:2009zz ; Carisch:2022ufd . The entanglement structure of mixed states is more complex, and there may be different ways to decompose correlations in a given mixed state density matrix. It is challenging to distinguish classical correlations from quantum entanglement in statistical mixtures. Several quantities have been proposed to measure entanglement in mixed states, such as entanglement of purification Takayanagi:2017knl ; Nguyen:2017yqw , reflected entropy Dutta:2019gen , odd entanglement entropy Tamaoka:2018ned , and entanglement negativity Vidal:2002zz ; Calabrese:2012ew ; Kudler-Flam:2018qjo .
Our primary focus is on using the quantity for entanglement in mixed states called reflected entropy Dutta:2019gen ; Jeong:2019xdr ; Akers:2021pvd ; Akers:2022max in the AdS3/CFT2 setup when the central charge $c$ is large Hartman:2013mia ; Fitzpatrick:2014vua ; Perlmutter:2015iya ; Anous:2016kss . Through this approach, we aim to extend the scope of calculations and strengthen the validity of AdS/CFT correspondence. We revisit and integrate previous studies, expanding calculations as far as possible. Specifically, the reason why we focus on reflected entropy is that it exhibits a different structure compared to entanglement entropy. Both entanglement entropy and reflected entropy can be expressed in terms of correlation functions in conformal field theory (CFT) Calabrese:2004eu ; Calabrese:2009qy . However, from the perspectives of $1/c$ expansion and conformal blocks, reflected entropy offers a wider range of scenarios.
@@
The gravity dual of reflected entropy is known as the entanglement wedge cross section in AdS spacetime Takayanagi:2017knl ; Nguyen:2017yqw . When we consider the entanglement wedge Czech:2012bh ; Wall:2012uf ; Headrick:2014cta associated with two subregions on the CFT, only one non-trivial phase of the entanglement wedge cross section emerges. However, in the context of BCFT, the dual geometry incorporates the end of the world brane Karch:2000gx ; Takayanagi:2011zk ; Fujita:2011fp ; Nozaki:2012qd . The presence of the end of the world brane allows the entanglement wedge associated with two subregions to involve a larger set of phases of the entanglement wedge cross section. For recent developments associated with reflected entropy and entanglement wedge cross section in AdS/BCFT, see Kusuki:2022bic ; Afrasiar:2022fid ; Afrasiar:2023jrj ; Basu:2023jtf ; Basak:2023bnc .
@@
This paper is organized as follows. In Section 2, we begin with the BCFT calculation. In Section 2.1, we introduce the assumptions of our calculations and review the reflected entropy in BCFT employing the doubling trick. We then calculate the $m-1$ correction to the reflected entropy in Section 2.2. Moving into the gravity calculation in Section 3, we review the entanglement wedge cross section without backreaction in Section 3.1. By using the conformal transformation, we calculate the entanglement wedge cross section involving backreaction from the cosmic brane in Section 3.2. To extend the scope of the calculation, we calculate the entanglement wedge cross section in a more general configuration of boundary intervals in Section 4.1 and show the consistency with the $m-1$ corrections in the AdS/CFT setup in Section 4.2. We conclude this paper by summarizing our results and outlook in Section 5.
@@ 2.1 Reflected entropy in BCFT2
which can be interpreted as the entanglement entropy of the reduced density matrix $\rho_{AA^{\star}}$. Reflected entropy is a well-defined quantity which includes both quantum and classical correlations between $A$ and $B$.
@@
When the size of the interval $A$ is smaller than that of the interval $B$, the numerator in equation (17) is factorized by
@@
Since the one-point function at $x_{2}$ is canceled by the denominator after taking the limit as $n$ approaches 1, the reflected entropy is determined by the one-point function at $x_{1}$. By using the doubling trick, the one-point function at $x_{1}$ is computed by the two-point function as shown in Figure 1.
@@
When the size of the interval $A$ is larger than that of the interval $B$, the numerator in equation (17) can be evaluated by
@@
where we used the doubling trick. To be precise, we should consider the contributions from equation (23) and another three-point function $\langle\sigma_{g_{B}}(-x_{2})\sigma_{g_{A}g_{B}^{-1}}(-x_{1})\sigma_{g_{B}^{-1}}(x_{2})\rangle$ in the chiral CFT (See Figure 2). Instead of analyzing them, we evaluate both holomorphic and antiholomorphic parts of equation (23). Consequently, the reflected entropy is determined as follows
@@
which is illustrated in Figure 2. The three-point function in the numerator is computed as
@@ 2.2 $m-1$ correction to reflected entropy in BCFT2
where we defined the rescaled conformal dimensions $\epsilon_{i}$ for each twist operators in equation (31) on the $mn$ replica sheets. Since we consider holographic CFTs, the identity block with $\epsilon_{p}=0$ becomes the dominant contribution in the conformal block expansion. In this case, $b_{p}$ is simply given by $b_{p}=1$. Consequently, for the partition function $Z_{nm}$ in equation (2.2) in the holomorphic part, we obtain
@@
The leading order term in equation (41) aligns well with the reflected entropy for phase I (22).
@@
In phase II, the numerator of the Renyi reflected entropy (2.2) can be evaluated using equation (23). The Renyi reflected entropy for phase II is then expressed as
@@
The leading order term in equation (46) reproduces the reflected entropy for phase II (27).
@@ 3.1 Entanglement wedge cross section in AdS3/BCFT2
As depicted in Figure 3, this phase of the entanglement wedge cross section is anchored to the end of the world brane. In this configuration, the entanglement wedge cross section is determined by minimizing the area of a surface that begins at $x_{1}$ and extends to a point on the end of the world brane within Poincare geometry. The calculation for this cross section is expressed as
@@ 4.1 General adjacent intervals
For the adjacent two intervals that are not attached to the boundary of BCFT, there are two possible configurations of the Ryu-Takayanagi (RT) surface of the subregion $A\cup B$: the connected configuration and the disconnected configuration. In the disconnected configuration, each endpoint of the subregion $A\cup B$ is anchored to the end of the world brane, creating two disconnected RT surfaces (Figure 4). In the connected configuration, two end points of the subregion $A\cup B$ are connected through a single RT surface (Figure 5).
@@
Here, for the sake of computational convenience, the cross ratios are defined in three cases, but the final results are written in terms of two cross ratios, $z_{+}$ and $z_{-}$. Additionally, we consider the case where the end of the world brane is orthogonal to the boundary of AdS spacetime, allowing us to ignore its backreaction effects. That is, we set $S_{\mathrm{bdy}}=0$ and $\theta=0$ in Figure  4.
@@
For the disconnected configuration, there are two disconnected RT surfaces that end at the end of the world brane. One of them starts at the point $x_{1}$ on the boundary and the other one starts at the point $x_{3}$ at the boundary. In this case, there are three possible phases of the entanglement wedge cross section (See Figure 4). For any phase, the entanglement wedge cross section extends from the point $x_{2}$ on the boundary of AdS spacetime. Depending on how the other end of the entanglement wedge cross section is located, we have three different phases as follows:
@@
This gravity calculation, along with Figure 4, is useful for understanding the corresponding computation in the CFT framework. The reflected entropy on the CFT side involves a six-point function in its numerator. At leading order, the calculation is related to determining the two-point function since the six-point function does factorize for $m=1$. However, if we consider the $m-1$ correction, it does not factorize. On the gravity side, the origin of this phenomenon can be understood by the influence of the cosmic brane anchored at the boundary points $x_{1}$ or $x_{3}$. In the phase I diagram in Figure 4, while each cosmic brane affects the entanglement wedge, their influences do not overlap at linear order; instead, the overlap is captured by higher-order corrections. As a result, the reflected entropy, which is related to a six-point function on the CFT side, would be expressed in terms of two four-point functions. This is evident from the expression of $C_{1}(z_{+},z_{-})$, where it is written in terms of two cross ratios, but the $z_{+}$ and $z_{-}$ function parts are completely separated.
@@
Unfortunately, we do not have results for the Rényi reflected entropy for two intervals, $A=[x_{1},x_{2}]$ and $B=[x_{2},x_{3}]$, computed in the BCFT setup. One interesting aspect of our results for the entanglement wedge cross section is that $C_{2}(z_{+},z_{-})$ may be related to an $\epsilon$-expansion of five-point functions in the CFT. As can be seen in the phase II diagram in Figure 4, the backreaction of the cosmic brane affects both the other cosmic brane and the entanglement wedge cross section, leading to the expectation that the linear order in $m-1$ is associated with a non-factorized five-point function. The expression for $C_{2}(z_{+},z_{-})$ is given in terms of two cross ratios, $z_{+}$ and $z_{-}$, and is intricately entangled, unlike in the case of phase I, where the components are separated.
@@
For the connected configuration, $x_{1}$ and $x_{3}$ are two boundary endpoints of the RT surface. In this case, there is only one phase of the entanglement wedge cross section that starts at the point $x_{2}$ on the boundary and ends at the RT surface (See Figure 5).
@@
As discussed, for the connected RT configuration of general adjacent intervals in AdS/BCFT setup, the entanglement wedge cross section is equivalent to that of AdS/CFT setup since the end of the world brane does not contribute. Therefore, (87) represents the entanglement wedge cross section involving the $m-1$ correction to for the connected RT configuration in AdS/BCFT. In order to match the zeroth order result in (108), $\delta$ should equal the UV cutoff $\delta_{x}$. In the next section, using the same strategy we will reproduce a result from phase II of adjacent intervals anchored to the boundary of BCFT, as derived in Section 3.2.

### v2

@@ 1 Introduction
There are various physical quantities used to measure the entanglement of quantum systems. Among them, entanglement entropy is a well-defined entanglement measure for pure states. However, it is not suitable as a measure for the entanglement of mixed states Horodecki:2009zz ; Carisch:2022ufd . The entanglement structure of mixed states is more complex, and there may be different ways to decompose correlations in a given mixed state density matrix. It is challenging to distinguish classical correlations from quantum entanglement in statistical mixtures. Several quantities have been proposed to study entanglement in mixed states, such as entanglement of purification Takayanagi:2017knl ; Nguyen:2017yqw , reflected entropy Dutta:2019gen , odd entanglement entropy Tamaoka:2018ned , and entanglement negativity Vidal:2002zz ; Plenio:2005cwa ; Calabrese:2012ew ; Kudler-Flam:2018qjo .
Our primary interest lies in reflected entropy Dutta:2019gen ; Jeong:2019xdr ; Akers:2021pvd ; Akers:2022max , a quantity designed to probe entanglement in mixed states. Although it is not a correlation measure for a generic state Hayden:2023yij , reflected entropy provides valuable insight into mixed state entanglement, particularly in the context of tripartite systems  Hayden:2021gno ; Akers:2019gcv . In this work, we investigate reflected entropy within the AdS3/CFT2 framework at large central charge $c$ Hartman:2013mia ; Fitzpatrick:2014vua ; Perlmutter:2015iya ; Anous:2016kss , where its monotonicity has been established for certain classes of states in holographic field theories Dutta:2019gen . Through this approach, we aim to extend the scope of calculations and strengthen the validity of AdS/CFT correspondence. We revisit and integrate previous studies, expanding calculations as far as possible. Specifically, the reason why we focus on reflected entropy is that it exhibits a different structure compared to entanglement entropy. Both entanglement entropy and reflected entropy can be expressed in terms of correlation functions in conformal field theory (CFT) Calabrese:2004eu ; Calabrese:2009qy . However, from the perspectives of $1/c$ expansion and conformal blocks, reflected entropy offers a wider range of scenarios.
@@
The gravity dual of reflected entropy is known as the entanglement wedge cross section in AdS spacetime Takayanagi:2017knl ; Nguyen:2017yqw . When we consider the entanglement wedge Czech:2012bh ; Wall:2012uf ; Headrick:2014cta associated with two subregions on the CFT, only one non-trivial phase of the entanglement wedge cross section emerges. However, in the context of BCFT, the dual geometry incorporates the end of the world brane Karch:2000gx ; Takayanagi:2011zk ; Fujita:2011fp ; Nozaki:2012qd ; Geng:2021iyq ; Geng:2022dua ; Geng:2024xpj . The presence of the end of the world brane allows the entanglement wedge associated with two subregions to involve a larger set of phases of the entanglement wedge cross section. For recent developments associated with reflected entropy and entanglement wedge cross section in AdS/BCFT and similar settings, see Sahraei:2021wqn ; Chen:2022fte ; Kusuki:2022bic ; Vasli:2022kfu ; Afrasiar:2022fid ; Afrasiar:2023jrj ; Basu:2023wmv ; Basu:2023jtf ; Lin:2023ajt ; Basak:2023bnc .
@@
This paper is organized as follows. In Sec. 2, we begin with the BCFT calculation. In Sec. 2.1, we introduce the assumptions of our calculations and review the reflected entropy in BCFT employing the doubling trick. We then calculate the $m-1$ correction to the reflected entropy in Sec. 2.2. Moving into the gravity calculation in Sec. 3, we review the entanglement wedge cross section without backreaction in Sec. 3.1. By using the conformal transformation, we calculate the entanglement wedge cross section involving backreaction from the cosmic brane in Sec. 3.2. To extend the scope of the calculation, we calculate the entanglement wedge cross section in a more general configuration of boundary intervals in Sec. 4.1 and show the consistency with the $m-1$ corrections in the AdS/CFT setup in Sec. 4.2. We conclude this paper by summarizing our results and outlook in Sec. 5.
@@ 2.1 Reflected entropy in BCFT2
which can be interpreted as the entanglement entropy of the reduced density matrix $\rho_{AA^{\star}}$. Reflected entropy is a useful quantity for analyzing entanglement in mixed states, which reflects both quantum and classical correlations between $A$ and $B$, though the reflected entropy of some states violates a condition for a measure of correlations.
@@
When the size of the interval $A$ is much smaller than that of the interval $B$, the four-point vacuum conformal block in the OPE channel for two pairs, $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})\sigma_{g_{B}g_{A}^{-1}}(x_{1})$ and $\sigma_{g_{B}}(-x_{2})\sigma_{g_{B}^{-1}}(x_{2})$, may be dominant for the numerator in equation (19). In addition, if the OPE for one of the pairs is simplified by an assumption that the twist operator can be regarded as the identity operator at $m,n\rightarrow 1$, the four-point function may be represented by lower-point functions. Under the above assumptions, we treat $\sigma_{g_{B}}(-x_{2})$ and $\sigma_{g_{B}^{-1}}(x_{2})$ as the identity operators at $m\to 1$ and decouple it from $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})\sigma_{g_{B}g_{A}^{-1}}(x_{1})$. [^4] If we treat both $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})$ and $\sigma_{g_{B}g_{A}^{-1}}(x_{1})$ as the identity operators at $n\to 1$, we cannot obtain a nonzero result. Strictly speaking, one must pay attention to the order of the limits $m\to 1$ and $n\to 1$ Kusuki:2019evw , but we are not concerned with that in this calculation. At least, our CFT calculation is consistent with the gravity computation in the next section. Confirming the assumption of the dominant channel is a future work. Then, the numerator in equation (17) is factorized by
@@
Since the one-point function at $x_{2}$ is canceled by the denominator after taking the limit as $n$ approaches 1, the reflected entropy is determined by the one-point function at $x_{1}$. By using the doubling trick, the one-point function at $x_{1}$ is computed by the two-point function as shown in Fig. 1.
@@
When the size of the interval $A$ is much larger than that of the interval $B$, the four-point conformal block in the OPE channel for two pairs, $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})\sigma_{g_{B}}(-x_{2})$ and $\sigma_{g_{B}g_{A}^{-1}}(x_{1})\sigma_{g_{B}^{-1}}(x_{2})$, may be dominant for the numerator in equation (19). In phase II, we assume that either $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})$ or $\sigma_{g_{B}g_{A}^{-1}}(x_{1})$ can be treated as the identity operator at $n\to 1$. For example, if we treat $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})$ as the identity operator, the numerator in equation (17) can be evaluated by
@@
where we used the doubling trick. To be precise, we should consider the contributions from equation (23) and another three-point function $\langle\sigma_{g_{B}}(-x_{2})\sigma_{g_{A}g_{B}^{-1}}(-x_{1})\sigma_{g_{B}^{-1}}(x_{2})\rangle$ in the chiral CFT (See Fig. 2). Instead of analyzing them, we evaluate both holomorphic and antiholomorphic parts of equation (23) for the full correlator, not the chiral part. Thanks to the symmetric configurations, our computation of the full correlator (23) matches with the chiral parts of the two configurations in Fig. 2. Consequently, the reflected entropy is determined as follows
@@
which is illustrated in Fig. 2. The three-point function in the numerator is computed as
@@ 2.2 $m-1$ correction to reflected entropy in BCFT2
where we defined the rescaled conformal dimensions $\epsilon_{i}$ for each twist operators in equation (31) on the $mn$ replica sheets. Since we consider holographic CFTs, the identity block with $\epsilon_{p}=0$ becomes the dominant contribution in the conformal block expansion. [^6] In our computation, we assume that the Virasoro identity block is dominant. In an orbifold CFT, it would be better to calculate using the orbifold block instead of the Virasoro block. Our CFT computation using the Virasoro block agrees one of the saddle solutions on the gravity side as shown in the next section. If we use the orbifold block instead of the Virasoro block, the CFT calculation in that case might be consistent with another saddle solution on the gravity side. It is an important future work to confirm this. In this case, $b_{p}$ is simply given by $b_{p}=1$. Consequently, for the partition function $Z_{nm}$ in equation (2.2) in the holomorphic part, we obtain
@@
The leading order term in equation (41) aligns well with the reflected entropy for phase I, as given in (22). The results in equations (40) and (41) demonstrate how the $m-1$ correction to the reflected entropy is captured through the operator product expansion involving the semi-classical Virasoro block. In particular, the second term in equation (40) explicitly represents the non-perturbative dependence of the Renyi reflected entropy on $m$, while its expansion in equation (41) reveals the leading order correction in small $m-1$ limit. It is important to note that the Renyi reflected entropy depends explicitly on the cross-ratio $z$. As will be shown in the analysis of phase II, this $z$-dependence gives rise to a phase transition in the Renyi reflected entropy.
@@
In phase II, as in the discussion at the previous subsection, we also assume that the numerator of the Renyi reflected entropy (2.2) can be evaluated using equation (23) even for $m\neq 1$ because the value of $m$ is not important for treating either $\sigma_{g_{A}g_{B}^{-1}}(-x_{1})$ or $\sigma_{g_{B}g_{A}^{-1}}(x_{1})$ as the identity operator at $n\to 1$. The Renyi reflected entropy for phase II is then expressed as
@@
The leading order term in equation (46) reproduces the reflected entropy for phase II as given in (27). Equations (45) and (46) illustrate how the $m-1$ correction to the reflected entropy appears in phase II. The $\log m$ term in (45), originating from the OPE coefficient in (26), captures the non-perturbative $m$-dependence. The Renyi reflected entropy at a given cross-ratio $z$ is determined by the minimum of equations (40) and (45). This behavior demonstrates how the presence of the boundary in AdS/BCFT modifies the entanglement structure, resulting in a phase transition in the Renyi reflected entropy—a boundary-induced effect that is absent in the standard AdS/CFT case.
@@ 3.1 Entanglement wedge cross section in AdS3/BCFT2
As depicted in Fig. 3, this phase of the entanglement wedge cross section is anchored to the end of the world brane. In this configuration, the entanglement wedge cross section is determined by minimizing the area of a surface that begins at $x_{1}$ and extends to a point on the end of the world brane within Poincare geometry. The calculation for this cross section is expressed as
@@ 3.2 Backreaction on entanglement wedge cross section
Equation (64) illustrates how the $m-1$ correction to the entanglement wedge cross section captures the effect of cosmic brane backreaction in the bulk, specifically when the entanglement wedge terminates on the end-of-the-world brane. The second term, which explicitly depends on $m$, arises from the deformation of the bulk geometry induced by the brane tension. To express this correction in terms of the cosmic brane tension $T_{m}=\frac{m-1}{4mG_{N}}$, we rewrite the second term in (64) as
$\frac{1}{4G_{N}}\log\left[\frac{z^{2G_{N}T_{m}}-z^{1-2G_{N}T_{m}}}{(1-z)(1-4G_{N}T_{m})}\right].$ (66)
In the limit $T_{m}\to 0$, this correction vanishes, and we recover the entanglement wedge cross section in the absence of backreaction. Since $0\leq T_{m}<\frac{1}{4G_{N}}$ for $m\in[1,\infty)$, the correction term is always negative within this range, indicating that the backreaction of the cosmic brane reduces the area of the entanglement wedge cross section.
@@
Equation (68) captures the $m-1$ correction to the entanglement wedge cross section due to the backreaction of a cosmic brane with tension $T_{m}=\frac{m-1}{4mG_{N}}$. The $\log m$ term reflects the conical defect induced by the brane, and can be rewritten as $\log(1-4G_{N}T_{m})^{-1}$, making its dependence on the brane tension explicit. As $T_{m}\to 0$, the correction vanishes, and we recover the result without backreaction. In contrast, as $m\to\infty$, the brane tension approaches $\frac{1}{4G_{N}}$, causing the correction term—and hence the entanglement wedge cross section—to diverge. The reason for this divergence is that the place where the cosmic brane at $m\to\infty$ is located corresponds to the horizon at $r=r_{h}=\frac{1}{m}\to 0$ in the metric (54). In phase II, the entanglement wedge cross section is anchored to the cosmic brane, leading to a divergent area because the distance to the horizon at $r=r_{h}=\frac{1}{m}\to 0$ is divergent. In contrast, in phase I, the entanglement wedge cross section does not touch the cosmic brane as seen in Fig. 3, so the area remains finite.
Importantly, the consistency between the entanglement wedge cross sections in equations (64) and (68) and the Renyi reflected entropies in equations (40) and (45) further solidifies the duality between boundary CFT calculations and bulk gravitational descriptions. This confirms that the $m-1$ expansion in the CFT precisely encodes the gravitational backreaction in the bulk.
@@ 4.1 General adjacent intervals
For the adjacent two intervals that are not attached to the boundary of BCFT, there are two possible configurations of the Ryu-Takayanagi (RT) surface of the subregion $A\cup B$: the connected configuration and the disconnected configuration. In the disconnected configuration, each endpoint of the subregion $A\cup B$ is anchored to the end of the world brane, creating two disconnected RT surfaces (Fig. 4). In the connected configuration, two end points of the subregion $A\cup B$ are connected through a single RT surface (Fig. 5).
@@
Here, for the sake of computational convenience, the cross ratios are defined in three cases, but the final results are written in terms of two cross ratios, $z_{+}$ and $z_{-}$. Additionally, we consider the case where the end of the world brane is orthogonal to the boundary of AdS spacetime, allowing us to ignore its backreaction effects. That is, we set $S_{\mathrm{bdy}}=0$ and $\theta=0$ in Fig. 4.
@@
For the disconnected configuration, there are two disconnected RT surfaces that end at the end of the world brane. One of them starts at the point $x_{1}$ on the boundary and the other one starts at the point $x_{3}$ at the boundary. In this case, there are three possible phases of the entanglement wedge cross section (See Fig. 4). For any phase, the entanglement wedge cross section extends from the point $x_{2}$ on the boundary of AdS spacetime. Depending on how the other end of the entanglement wedge cross section is located, we have three different phases as follows:
@@
This gravity calculation, along with Fig. 4, is useful for understanding the corresponding computation in the CFT framework. The reflected entropy on the CFT side involves a six-point function in its numerator. At leading order, the calculation is related to determining the two-point function since the six-point function does factorize for $m=1$. However, if we consider the $m-1$ correction, it does not factorize. On the gravity side, the origin of this phenomenon can be understood by the influence of the cosmic brane anchored at the boundary points $x_{1}$ or $x_{3}$. In the phase I diagram in Fig. 4, while each cosmic brane affects the entanglement wedge, their influences do not overlap at linear order; instead, the overlap is captured by higher-order corrections. As a result, the reflected entropy, which is related to a six-point function on the CFT side, would be expressed in terms of two four-point functions. This is evident from the expression of $C_{1}(z_{+},z_{-})$, where it is written in terms of two cross ratios, but the $z_{+}$ and $z_{-}$ function parts are completely separated.
This result provides a holographic framework for analyzing the $\epsilon$-expansion of conformal blocks in six-point correlation functions, for the operator configuration corresponding to this phase. In particular, the calculation corresponds to the $m-1$ correction to the following six-point correlation function in the CFT [^11] The relation between the reflected entropy and the correlation function is discussed in A.1.
$\left\langle\sigma_{g_{B}}(-x_{3})\sigma_{g_{B}^{-1}g_{A}}(-x_{2})\sigma_{g_{A}^{-1}}(-x_{1})\sigma_{g_{A}}(x_{1})\sigma_{g_{B}g_{A}^{-1}}(x_{2})\sigma_{g_{B}^{-1}}(x_{3})\right\rangle_{\mathrm{CFT}}\,.$ (75)
This correlator consists of four heavy operators and two light operators in the case where $n=1$ and $m\neq 1$. The separation of contributions from $z_{+}$ and $z_{-}$ in equation (73) suggests that the $m-1$ correction to the six-point function in the CFT can be effectively captured by the $\epsilon$-expansion of four-point conformal blocks.
@@
Unfortunately, we do not have results for the Renyi reflected entropy for two intervals, $A=[x_{1},x_{2}]$ and $B=[x_{2},x_{3}]$, computed in the BCFT setup. One interesting aspect of our results for the entanglement wedge cross section is that $C_{2}(z_{+},z_{-})$ may be related to an $\epsilon$-expansion of five-point functions in the CFT. As can be seen in the phase II diagram in Fig. 4, the backreaction of the cosmic brane affects both the other cosmic brane and the entanglement wedge cross section, leading to the expectation that the linear order in $m-1$ is associated with a non-factorized five-point function. The expression for $C_{2}(z_{+},z_{-})$ is given in terms of two cross ratios, $z_{+}$ and $z_{-}$, and is intricately entangled, unlike in the case of phase I, where the components are separated.
This result suggests that the $m-1$ correction to the entanglement wedge cross section in phase II provides a holographic perspective on the $\epsilon$-expansion of the five-point function in CFT, specifically
$\left\langle\sigma_{g_{B}}(-x_{3})\sigma_{g_{A}^{-1}}(-x_{1})\sigma_{g_{A}}(x_{1})\sigma_{g_{B}g_{A}^{-1}}(x_{2})\sigma_{g_{B}^{-1}}(x_{3})\right\rangle_{\mathrm{CFT}}\,,$ (83)
which consists of four heavy operators and one light operator in the limit $n=1$ and $m\neq 1$. In conjunction with the relation between equations (73) and (75), this observation indicates that gravity calculations may provide novel insights into the structure of conformal blocks within the holographic framework.
@@
For the connected configuration, $x_{1}$ and $x_{3}$ are two boundary endpoints of the RT surface. In this case, there is only one phase of the entanglement wedge cross section that starts at the point $x_{2}$ on the boundary and ends at the RT surface (See Fig. 5).
@@
As discussed, for the connected RT configuration of general adjacent intervals in AdS/BCFT setup, the entanglement wedge cross section is equivalent to that of AdS/CFT setup since the end of the world brane does not contribute. Therefore, (90) represents the entanglement wedge cross section involving the $m-1$ correction to for the connected RT configuration in AdS/BCFT. In order to match the zeroth order result in (111), $\delta$ should equal the UV cutoff $\delta_{x}$. In the next section, using the same strategy we will reproduce a result from phase II of adjacent intervals anchored to the boundary of BCFT, as derived in Sec. 3.2.